OpenStudy (anonymous):

x/x+3=3/19 Find the LCD for denominator, then solve. Please show all your work, THANK YOU!

OpenStudy (anonymous):

seems pretty pointless to find common denominators , when you could just cross multiply

OpenStudy (anonymous):

idk haha thats just what the book says to do, or how to do it, but you have a good point.

OpenStudy (anonymous):

19x = 3(x+3) 19x =3x +9 16x =9 x= 9/16

OpenStudy (anonymous):

just cross multiply , the textbook is asking a stupi question